OpenStudy (anonymous):

Find the values of x in the following equation:|-5x + -2| + 3 = 11 A. x = -3.2 or x = -2 B. x = -1.2 or x = 2 C. x = -2 or x = 1.2 D. x = -2 or x = 3.2

Directrix (directrix):

|-5x + -2| + 3 = 11 |-5x + -2| = 8 Agree @AriannaA

OpenStudy (anonymous):

May I ask how you got that? :o

Directrix (directrix):

|-5x + -2| + 3 = 11 |-5x + -2| + 3 + (-3) = 11 + (-3)

Directrix (directrix):

|-5x + -2| + 3 + (-3) = 11 + (-3) |-5x + -2| + 0 = 8

Directrix (directrix):

|-5x + -2| + 0 = 8 |-5x + -2| = 8

Directrix (directrix):

Is that okay?

OpenStudy (anonymous):

So what does X equal? D: aslkdjfsddsf

Directrix (directrix):

We are moving by steps here to get in a position to apply the definition of the absolute value function to solve the problem.

Directrix (directrix):

| x | = x if x is positive or zero, but -x if x is negative

Directrix (directrix):

By definition of absolute value: -5x + -2 = 8 OR -5x + -2 = -8 Solve each of the two equations for x. @AriannaA

OpenStudy (anonymous):

I don't understand. >.<

Directrix (directrix):

It is a good idea to memorize the definition of absolute value. | x | = x if x is positive or zero, but -x if x is negative

Directrix (directrix):

|8 |= 8 Also, |-8| = 8

Directrix (directrix):

So, what is between the bars is equal to 8 or to -8 because both 8 and -8 have the same absolute value.

Directrix (directrix):

-5x + -2 is between the absolute value bars.

Directrix (directrix):

So, -5x + -2 = 8 OR -5x + -2 = -8 ----------- Solve -5x + -2 = 8 for x. Then, solve -5x + -2 = -8 for x. The question has two answers.

Directrix (directrix):

@AriannaA Please check to see if the options are correctly posted. Thanks.